In a genetics class, 6

 students have GREEN eyes, 3

 students have BLUE eyes, and 11

 have HAZEL eyes. If a single student is picked at random, what is the probability that their eyes are GREEN or HAZEL?


1
Expert's answer
2021-10-20T18:02:42-0400

probability of green eyes:

"p_g=6\/(6+3+11)=3\/10"


probability of hazel eyes:

"p_h=11\/(6+3+11)=11\/20"


probability of green or hazel eyes:

"p=3\/10+11\/20=17\/20"